OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (–3, 7); Slope: 4

pooja195 (pooja195):

\[y-y1=m(x-x1)\]

pooja195 (pooja195):

now just plug in the numbers do you know how?

OpenStudy (anonymous):

I have absolutely no idea how to do this.

pooja195 (pooja195):

ok i just gave you the formula now we plug in the numbers (x,y) what is the y in our equation what number?

OpenStudy (anonymous):

7?

pooja195 (pooja195):

\[y-7=m(x-x1)\] m=slope so whats the slope?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

y – 7 = 4(x – 3)?

pooja195 (pooja195):

\[y-7=4(x-x1)\] whats the x

OpenStudy (anonymous):

-12?

pooja195 (pooja195):

\[y – 7 = 4(x –(- 3))\]

OpenStudy (anonymous):

i cant see that

pooja195 (pooja195):

refresh your page